# Themed Issue in Origami and Kirigami Mechanics

### v1

@@ Abstract
Yoshimura origami is a classical folding pattern that has inspired many deployable structure designs. Its applications span from space exploration, kinetic architectures, and soft robots to even everyday household items. However, despite its wide usage, Yoshimura has been fixated on a set of design constraints to ensure its flat-foldability. Through extensive kinematic analysis and prototype tests, this study presents a new Yoshimura that intentionally defies these constraints. Remarkably, one can impart a unique meta-stability by using the Golden Ratio angle to define the triangular facets of a generalized Yoshimura. As a result, when its facets are strategically popped out, a “Golden Ratio Yoshimura” boom with $m$ modules can be theoretically reconfigured into $8^{m}$ geometrically unique and load-bearing shapes. This result not only challenges the existing design norms but also opens up a new avenue to create deployable and versatile structural systems.
@@ 1 Introduction
Deployable structures have garnered considerable interest over the past decades due to their unique ability to transform from a compact volume to a larger-scale, load-bearing shape [fenci2017, friedman2011a, zhang2021]. Their applications are diverse. For example, they have been the essential components for satellites, exoplanet rovers, and space telescopes [puig2010, santiago2013]. They will continue to play a vital role in the future in in-orbit fabrication [liu2022a, ekblaw2019] and space habit construction [yang2023, liu2022b, hill2010]. Here on Earth, deployable structures are ideal for making military and civilian shelters [thrall2014, mira2014] and are even used to develop novel soft robots [troise2021, sedal2020, blumenschein2018, kaufmann2022].
@@
Figure 1: A big-picture overview of the Golden Ratio Yoshimura as a deployable and reconfigurable structure. (a) A 3D-printed, meter-scale prototype can be easily configured into different meta-stable shapes for different purposes, including a fully folded (compact) configuration, a fully deployed configuration, a load-bearing arc, or letters “VT” and “CU,” representing the institute affiliation of the authors. Notice the 3D printed sample weighs 225.6 grams, and a second Yoshimura was used to finish the letter “T.” (b). The buckling pattern of a cylindrical thin shell reveals the origin of Yoshimura. Image adapted from [singer1982status] with permission.
@@
Despite such wide usage, deployable Yoshimura has been fixated on a design rule in the current literature: The shape of its triangular facets must be constrained by the number of rhombus creases along the cylinder’s circumference. This rule can ensure the flat-foldability of Yoshimura. However, we showed that by intentionally breaking these established design rules, one can significantly enrich the kinematic configuration space of Yoshimura and generate a unique “meta-stability.” In this way, one can selectively pop out pairs of triangular facets to settle the new Yoshimura into many different elastically stable and load-bearing configurations. Remarkably, and for the first time, we mathematically proved that, for a Yoshimura pattern with three rhombi along its circumference (aka. $n=3$ in Fig. 2), the required triangular facet shape to attain such meta-stability follows the Golden Ratio (aka. $\cot\beta=\varphi\approx 1.618$ in Fig. 1, 2). We refer to this new origami design as the “Golden Ratio Yoshimura,” and it can become the foundation for those, as mentioned earlier, massively reconfigurable and deployable structures.
@@ 1 Pop-out Kinematics
To further correlate these two angles to the dihedral facet angle ($\theta$) and mountain crease incline angle $\beta$, we started by calculating the height of the original triangular facet ($w$ as in Fig. 2), in that:
@@
$\begin{split}O&\equiv[0,0,0],\quad(\text{Reference})\\ A&\equiv[0,L\cos\eta+\frac{L}{2}\cos\alpha,w],\\ B&\equiv[\frac{L}{2},w\cos\theta,w\sin\theta],\\ R&\equiv[L\sin\eta,L\cos\eta,0].\end{split}$ (6)
@@
$|A-B|^{2}=L^{2}=\Big{[}\frac{L}{2}-0\Big{]}^{2}+\Big{[}w\cos\theta-(L\cos\eta+\frac{L}{2}\cos\alpha)\Big{]}^{2}+\Big{[}w\sin\theta-w\Big{]}^{2}.$ (7)
@@
$32\tan^{2}\beta(1-\sin\theta)+4\cos^{2}\eta+\cos^{2}\alpha-2\tan\beta\cos\theta(2\cos\eta+\cos\alpha)=$ 3 . (8)
@@
$|R-B|^{2}=\left[\frac{L}{2\cos\beta}\right]^{2}=\Big{[}L\sin\eta-\frac{L}{2}\Big{]}^{2}+\Big{[}L\cos\eta-w\cos\theta\Big{]}^{2}+\Big{[}0-w\sin\theta\Big{]}^{2}.$ (9)
@@
Using Eq. (4-10), we could fully define the geometry of a 1 pop-out Yoshimura module. These equations are highly non-linear and transcendental, so they were solved numerically. Figure 3(c) summarizes the solution showing the correlation between the dihedral facet angle $\theta$ and the Yoshimura design parameter $\beta$. Surprisingly, we observe no solution when $\beta<31.72\degree$, which is directly related to the Golden Ratio. This discovery guided us to define the “Golden Ratio Yoshimura” pattern for which the dihedral angle ($\theta$) reduces to zero. As a result, the vertices B and C on the top interface triangle will coincide with their counterpart B’ and C’ on the base triangle (i.e., the top surface and base triangle share a common edge). For any $\beta<31.72\degree$, the two adjacent facets PCO and PC’O would interfere and create geometric frustration, resulting in unreliable mechanics.
@@
$E\equiv[0,w\cos\theta,w\;\sin\theta],$ (11)
@@
$\tan\frac{\gamma}{2}=\frac{w-w\sin\theta}{L\cos\eta+\frac{L}{2}\cos\alpha-w\cos\theta}.$ (12)
@@ 2 Pop-out Kinematics
A 2 pop-out configuration emerges when two valley creases in the mid-section of a module are pushed outwards. One can reach this configuration either by popping out an additional valley fold from a 1 pop-out state or popping in a rhombus from the 3 pop-out configuration. Figure 3(e) details the geometry of this configuration with three different views. Again, we assume the top and base interface triangles remained unchanged.
@@
$\begin{split}C&\equiv[0,-w\cos\theta,w\sin\theta],\\ B&\equiv[\frac{L}{2},\frac{L}{2},w].\end{split}$ (13)
@@
$|B-C|^{2}=L^{2}=\Big{[}\frac{L}{2}-0\Big{]}^{2}+\Big{[}\frac{L}{2}-(-w\cos\theta)\Big{]}^{2}+\Big{[}w-w\sin\theta\Big{]}^{2}.$ (14)
@@
Figure 3(f) summarizes the solution of Eq. (15). Again, we observe that the dihedral angle between facet PCO and PC’O drops to zero when the sector angle $\beta$ equals the golden ratio angle of 31.72°. Any smaller $\beta$ angle will give a negative $\theta$, meaning that the two facets intersect and are not physically possible.
Therefore, we mathematically proved that the Golden Ratio angle is the lower bound of the $\beta$ value for kinematically admissible pop-out operations ($n=3$).
@@
$D\equiv[0,\frac{L}{2},w]$ (16)
@@
$\tan\frac{\gamma}{2}=\frac{w-w\sin\theta}{\frac{L}{2}+w\cos\theta}$ (17)
@@ 3.1. Forward Kinematics Formulation
The corresponding transformation parameters for each module are listed in Table 1. Finally, by treating each module as an independent building block, we could determine the kinematic state of the entire structure, by applying a chain of transformations from the $i^{th}$ and $j^{th}$ modules ($i<j$):
@@ 3.2. Exploring and Accessing the Configuration Space of Yoshimura
The Golden Ratio Yoshimura boom offers a versatile way to construct complex shapes by combining different modules. We can adjust the length scaling and curvature to achieve intricate shapes overall. For example, we could reach the maximum increment in length ($l^{\prime}$) by stacking either the 111 module or a pair of 1 and 2 pop-out modules with the same phase $\psi$ (e.g., 001 and 110 according to Table 1). The corresponding increment in length is $1/\varphi$. On the other hand, the minimum increment in length ($l$) is achieved with 000 modules, and the value is given by $1/\varphi\sqrt{1-\varphi^{2}/3}$.
@@
Figure 4(d) demonstrates the possibilities of these combinations.
@@ 4.2 Load Bearing Capacity Tests
Figure 6(a) summarizes the longitudinal compression test results (Instron 6500 with 100N load cell). Overall, the 3D-printed Yoshimura samples exhibited remarkable durability, giving repeatable mechanical behaviors even after extensive usage. The dual-material Yoshimura sample withstood over 50 Newtons of compression force before beginning to buckle and transitioning from the deployed 3 pop-out state to the folded 0 pop-out state eventually (the sample itself weighs about 23 grams). Such buckling and folding occurred in sequence, so each small peak on the force-deformation curve indicates an inward buckling (or pop-in) of one triangular facet pair. The Yoshimura sample fully folded after being compressed by 62mm, and once it transitioned into the folded state, the Yoshimura sample behaved like a soft linear spring. In contrast, the pure TPU Yoshimura sample could carry only about 25 Newtons of compression force. These compression tests demonstrated that the nylon facets can significantly enhance the overall load-bearing, so we only used the dual material sample for the subsequent 3-point bending tests.
@@
Figure 6(c) summarizes the tested external moment-bending angle relationships, elucidating the bending stiffness of the Yoshimura structure under various configurations. Overall, the Yoshimura could carry significantly more bending load in their 2 pop-out and 3 pop-out configurations than the 0 pop-out state, validating its advantage as a load-carrying reconfigurable structure. The bending response is generally linear, except for the 2 pop-out results. When loaded on the valley crease, the 2 pop-out Yoshimura exhibits a close-to-linear bending response. However, when loaded on the vertex, the 2 pop-out Yoshimura showed a significantly nonlinear behavior with two distant “snap-through” events. These snap-through events occurred when the popped-out facets, under compression in this loading condition, eventually buckled and folded inwards. This behavior is similar to what was observed in the longitudinal compression test.

### v3

@@ Abstract
Yoshimura origami is a classical folding pattern that has inspired many deployable structure designs. Its applications span from space exploration, kinetic architectures, and soft robots to even everyday household items. However, despite its wide usage, Yoshimura has been fixated on a set of design constraints to ensure its flat-foldability. Through extensive kinematic analysis and prototype tests, this study presents a new Yoshimura that intentionally defies these constraints. Remarkably, one can impart a unique meta-stability by using the Golden Ratio angle ($\cot^{-1}1.618\approx 31.72\degree$) to define the triangular facets of a generalized-Yoshimura (with $n=3$, where $n$ is the number of rhombi shapes along its cylindrical circumference). As a result, when its facets are strategically popped out, a “Golden Ratio Yoshimura” boom with $m$ modules can be theoretically reconfigured into $8^{m}$ geometrically unique and load-bearing shapes. This result not only challenges the existing design norms but also opens up a new avenue to create deployable and versatile structural systems.
@@ 1 Introduction
Deployable structures have garnered considerable interest over the past decades due to their unique ability to transform from a compact volume to a larger-scale, load-bearing shape [fenci2017, friedman2011a, zhang2021]. Their applications are diverse. For example, they have been the essential components
for satellites, exoplanet rovers, and space telescopes [puig2010, santiago2013]. They will continue to play a vital role in the future for in-orbit fabrication [liu2022a, ekblaw2019] and space habit construction [yang2023, liu2022b, hill2010]. Here on Earth, deployable structures are ideal for making military and civilian shelters [thrall2014, mira2014] and are even used to develop novel soft robots [troise2021, sedal2020, blumenschein2018, kaufmann2022].
@@
Figure 1: A big-picture overview of the Golden Ratio Yoshimura as a deployable and reconfigurable structure. (a) A 3D-printed, meter-scale prototype can be easily configured into different meta-stable shapes for different purposes, including a fully folded (compact) configuration, a fully deployed configuration, a load-bearing arc, or letters “VT” and “CU.” Notice the 3D printed sample weighs 225.6 grams, and a second Yoshimura was used to finish the letter “T.” (b). The buckling pattern of a cylindrical thin shell reveals the origin of Yoshimura. Image adapted from [singer1982status] with permission.
@@
Despite such wide usage, deployable Yoshimura has been fixated on a design rule in the current literature: The shape of its triangular facets must be constrained by the number of rhombus creases along the cylinder’s circumference. This rule can ensure the flat-foldability of Yoshimura. However, we showed that by intentionally breaking these established design rules, one can significantly enrich the kinematic configuration space of Yoshimura and generate a unique “meta-stability.” In this way, one can selectively pop out pairs of triangular facets to settle the new Yoshimura into many different elastically stable and load-bearing configurations. Note that this feature sacrifices the flat-foldability, as the more generalized Yoshimura pattern has a finite height when folded. Remarkably, and for the first time, we mathematically proved that, for a Yoshimura pattern with three rhombi along its circumference (aka. $n=3$ in Fig. 2), the required triangular facet shape to attain such meta-stability follows the Golden Ratio (aka. $\cot\beta=\varphi\approx 1.618$ in Fig. 1, 2). We refer to this new origami design as the “Golden Ratio Yoshimura,” and it can become the foundation for those, as mentioned earlier, massively reconfigurable and deployable structures.
@@ 1 Pop-out Kinematics
Here, $\eta$ must lie in $[0,\frac{\pi}{3}]$ to ensure a real-valued solution for angle $\alpha$. To further correlate these two angles to the dihedral facet angle ($\theta$) and mountain crease incline angle $\beta$, we started by calculating the height of the original triangular facet ($w$ as in Fig. 2), in that:
@@
$\begin{split}O&\equiv\left[0,0,0\right],\quad(\text{Reference})\\ A&\equiv\left[0,L\cos\eta+\frac{L}{2}\cos\alpha,w\right],\\ B&\equiv\left[\frac{L}{2},w\cos\theta,w\sin\theta\right],\\ R&\equiv\left[L\sin\eta,L\cos\eta,0\right].\end{split}$ (6)
@@
$|A-B|^{2}=L^{2}=\left(\frac{L}{2}-0\right)^{2}+\left(w\cos\theta-\left(L\cos\eta+\frac{L}{2}\cos\alpha\right)\right)^{2}+\left(w\sin\theta-w\right)^{2}.$ (7)
@@
$32\tan^{2}\beta(1-\sin\theta)+\left(2\cos\eta+\cos\alpha\right)^{2}-2\tan\beta\cos\theta(2\cos\eta+\cos\alpha)=$ 3 . (8)
@@
$|R-B|^{2}=\left(\frac{L}{2\cos\beta}\right)^{2}=\left(L\sin\eta-\frac{L}{2}\right)^{2}+\left(L\cos\eta-w\cos\theta\right)^{2}+\left(0-w\sin\theta\right)^{2}.$ (9)
@@
Using Eq. (4-10), we could fully define the geometry of a 1 pop-out Yoshimura module. These equations are highly non-linear and transcendental, so they were solved numerically. Figure 3(c) summarizes the numerical solution showing the correlation between the dihedral facet angle $\theta$ and the Yoshimura design parameter $\beta$. It is worth noting that, in the 0 pop-out configuration as shown in Fig. 2, the dihedral angle $\theta$ is the same for all three pairs of triangular facets. However, in the 1 pop-out configuration, $\theta$ becomes $90\degree$ for the popped facets (aka. PAR and PA’R pair in Fig. 3(b)), while the dihedral angle $\theta$ of the other two facet pairs (OBR-OB’R and OCP-OC’P) are non-zero and related to $\beta$.
To understand the lower limit of sector angle $\beta$ that allows this kinematically admissible configuration, we obtain the closed-form solution of $\beta$ when the dihedral angle $\theta$ vanishes. By substituting zero dihedral angle condition ($\theta=0$), we observe from the top view of Fig. 3(b) that $\eta=\frac{\pi}{2}-2\beta$. Combining this observation with Eq. (4) implies that $\beta$ is constrained to the interval $\left[\frac{\pi}{12},\frac{\pi}{4}\right]$. Applying these conditions to Eq. (10) and using the identity $\cos\left(\frac{\pi}{2}-x\right)=\sin(x)$ yields:
$32\tan^{2}\beta+\left(2\sin(2\beta)+\cos\alpha\right)^{2}-2\tan\beta\left(2\sin(2\beta)+\cos\alpha\right)=$ 3 . (11)
We then substitute $\alpha$ from Eq. (4) into Eq. (11), leading to a polynomial in $\tan\beta$:
$\left(\tan\beta-\varphi^{-1}\right)\left(\tan\beta-\varphi\right)\left(5\tan^{4}\beta+4\sqrt{5}\tan^{3}\beta-2\tan^{2}\beta-4\sqrt{5}\tan\beta-7\right)=0.$ (12)
Here, $\varphi$ denotes the Golden Ratio, given by $\frac{1+\sqrt{5}}{2}$. Given the constraints on $\beta$, $\tan\beta$ must lie within the interval $[2-\sqrt{3},1]$. Consequently, the only valid solution for Equation (12) is $\tan\beta=\frac{1}{\varphi}$. This implies $\cot\beta=\varphi$, which leads to $\beta\approx 31.72^{\circ}$. Hence, no solution exists for $\beta<31.72^{\circ}$. This result is intrinsically linked to the Golden Ratio and guided us to define the "Golden Ratio Yoshimura" pattern, where the dihedral angle $\theta$ of the folded facet pairs reduces to zero. As a result, the vertices B and C on the top interface triangle will coincide with their counterpart B’ and C’ on the base triangle (i.e., the top surface and base triangle share a common edge). For any $\beta<31.72\degree$, the two adjacent facets PCO and PC’O would interfere and create geometric frustration, resulting in unreliable mechanics.
@@
$E\equiv\left[0,w\cos\theta,w\;\sin\theta\right],$ (13)
@@
$\tan\frac{\gamma}{2}=\frac{w-w\sin\theta}{L\cos\eta+\frac{L}{2}\cos\alpha-w\cos\theta}{\color[rgb]{1,0,0}=\frac{1-\sin{\theta}}{2\tan{\beta}\cos{\eta}+\tan{\beta}\cos\alpha-\cos\theta}}.$ (14)
@@ 2 Pop-out Kinematics
A 2 pop-out configuration emerges when two valley creases in the mid-section of a module are pushed outwards. One can reach this configuration either by popping out an additional valley fold from a 1 pop-out state or popping in a rhombus from the 3 pop-out configuration. Fig. 3(e) details the geometry of this configuration with three different views. Again, we assume the top and base interface triangles remained unchanged.
@@
$\begin{split}C&\equiv\left[0,-w\cos\theta,w\sin\theta\right],\\ B&\equiv\left[\frac{L}{2},\frac{L}{2},w\right].\end{split}$ (15)
@@
$|B-C|^{2}=L^{2}=\left(\frac{L}{2}-0\right)^{2}+\left(\frac{L}{2}-(-w\cos\theta)\right)^{2}+\left(w-w\sin\theta\right)^{2}.$ (16)
@@
The dihedral angle ($\theta$) between facet PCO and PC’O can drop to zero since any negative $\theta$ will cause self-intersection between the two facets and are not physically possible. Substituting $\theta=0$ into Eq. (17) yields a quadratic equation of $\tan{\beta}$. Upon solving it with a quadratic formula and choosing the solution with a positive value, we again obtain $\tan{\beta}=\frac{-1+\sqrt{5}}{2}=\varphi^{-1}$, or $\beta\approx 31.72^{\circ}$.
Therefore, we obtain the closed-form analytical solution that the Golden Ratio angle is the lower bound of the $\beta$ value for kinematically admissible pop-out operations ($n=3$).
@@
$D\equiv\left[0,\frac{L}{2},w\right]$ (18)
@@
$\tan\frac{\gamma}{2}=\frac{w-w\sin\theta}{\frac{L}{2}+w\cos\theta}{\color[rgb]{1,0,0}=\frac{1-\sin{\theta}}{\tan{\beta}+\cos{\theta}}}.$ (19)
@@ 3.1. Forward Kinematics Formulation
The corresponding transformation parameters for each module are listed in Table 1 (Detailed calculations of parameters are given in A). Finally, by treating each module as an independent building block, we could determine the kinematic state of the entire structure, by applying a chain of transformations from the $i^{th}$ and $j^{th}$ modules ($i<j$):
@@ 3.2. Exploring and Accessing the Configuration Space of Yoshimura
The Golden Ratio Yoshimura boom offers a versatile way to construct complex shapes by combining different modules. We can adjust the length scaling and curvature to achieve intricate shapes overall. For example, we could reach the maximum increment in length ($l^{\prime}$) by stacking either the 111 module or a pair of 1 and 2 pop-out modules with the same phase $\psi$ (e.g., 001 and 110 according to Table 1). The corresponding increment in length is $1/\varphi\approx 0.618$. On the other hand, the minimum increment in length ($l$) is achieved with 000 modules, and the value is given by $1/\varphi\sqrt{1-\varphi^{2}/3}\approx 0.2205$.
@@
In Fig. 4(d) the various combinations of module configurations are depicted. These simulations are made by triangulating nodes calculated using the kinematic model and then plotted using a 3D visualizer based on Python.
@@ 4.2 Load Bearing Capacity Tests
Fig. 6(a) summarizes the longitudinal compression test results (Instron 6500 with 100N load cell). Overall, the 3D-printed Yoshimura samples showed robust performance, exhibiting repeatable mechanical behaviors even after extensive usage. The dual-material Yoshimura sample withstood over 50 Newtons of compression force before beginning to buckle and transitioning from the deployed 3 pop-out state to the folded 0 pop-out state eventually (the sample itself weighs about 23 grams). Such buckling and folding occurred in sequence, so each small peak on the force-deformation curve indicates an inward buckling (or pop-in) of one triangular facet pair. The Yoshimura sample fully folded after being compressed by 62mm, and once it transitioned into the folded state, the Yoshimura sample behaved like a soft linear spring. In contrast, the pure TPU Yoshimura sample could carry only about 25 Newtons of compression force. These compression tests demonstrated that the nylon facets can significantly enhance the overall load-bearing, so we only used the dual material sample for the subsequent 3-point bending tests.
@@
Fig. 6(c) summarizes the tested external moment-bending angle relationships, elucidating the bending stiffness of the Yoshimura structure under various configurations. Overall, the Yoshimura could carry significantly more bending load in their 2 pop-out and 3 pop-out configurations than the 0 pop-out state, validating its advantage as a load-carrying reconfigurable structure. The bending response is generally linear, except for the 2 pop-out results. When loaded on the valley crease, the 2 pop-out Yoshimura exhibits a close-to-linear bending response. However, when loaded on the vertex, the 2 pop-out Yoshimura showed a significantly nonlinear behavior with two distant “snap-through” events. These snap-through events occurred when the popped-out facets, under compression in this loading condition, eventually buckled and folded inwards. This behavior is similar to what was observed in the longitudinal compression test.
@@ 5 Conlusion and Discusion
Appendix A Forward Kinematics Parameters
Phase Angles ($\psi$)
As explained in Section 3, configurations 000, 100, 011, and 111 are symmetric about the y-z reference plane; hence the phase angle is 0. On the other hand, the asymmetry of 001, 010, 101, and 110 is compensated by a rotation about the z-axis with an angle of $2\pi/n$. In this study, $n=3$, so that the phase angle is $32\pi/$ 3. Based on Fig. 4(a), one can see that the 010 and 101 configurations need a rotation by $32\pi/$ 3 to become symmetric about the y-z reference plane. Similarly, 001 and 100 configurations need a rotation of $34\pi/$ 3 or $-2\pi/3$.
Tilt Angles ($\gamma$)
The tilt angles for the 000 and 111 configurations are 0 as the top and bottom planes are parallel. Furthermore, both the 1 pop-out configurations (001, 010, 100) and the 2 pop-out configurations (101, 110, 011) are congruent in terms of tilt, resulting in equal magnitudes of tilt angles in these cases. The magnitudes are calculated using Eq. (14) for 1 pop-out and Eq. (19) for 2 pop-out.
For the 1 pop-out configuration, one can substitute $\theta=0$, $\beta=\cot^{-1}\varphi$, and the values of $\alpha$ and $\eta$ calculated in Section 2 ($\eta=\pi/2-2\beta$, $\alpha=\tan^{-1}2$) to simplifying Eq. (14), yielding:
${\color[rgb]{1,0,0}\tan\left(\frac{\gamma}{2}\right)=\frac{\varphi}{1+\varphi},}$ (31)
which further simplifies based on the trigonometric identity $\sin{x}=\frac{\tan x}{\sqrt{1+\tan^{2}x}}$:
${\color[rgb]{1,0,0}|(\gamma_{\text{gold}})_{1-p/o}|=2\sin^{-1}\left(\frac{1}{\sqrt{3}\varphi}\right)\approx 41.81^{\circ}.}$ (32)
Interestingly, substituting $\theta=0$ into Eq. (19) yields the exact same solution:
${\color[rgb]{1,0,0}|(\gamma_{\text{gold}})_{2-p/o}|=2\sin^{-1}\left(\frac{1}{\sqrt{3}\varphi}\right)\approx 41.81.}$ (33)
Since the symmetric configurations 100 and 011 are tilted opposite to each other, we assign positive values to the 1 pop-out configuration and negative values to the 2 pop-out configuration.
Slant Heights ($d$)
For the 0 and 3 pop-out configurations, the slant height is the same as the module’s total height.
The height at the 000 configurations can be calculated using the equation for $h$ (Eq. (2)) and substituting the $\theta$ solution from Eq. (3) corresponding to the Golden Ratio, i.e., substituting $\beta=\cot^{-1}\varphi$, so that:
${\color[rgb]{1,0,0}(d_{\text{gold}})_{0-p/o}=\frac{1}{\varphi}\sin\left(\cos^{-1}\left(\varphi\tan\left(\frac{\pi}{6}\right)\right)\right)=\frac{1}{\varphi}\sqrt{1-\frac{\varphi^{2}}{3}}\approx 0.2205.}$ (34)
For the 1 and 2-pop out configurations, the slant height is defined as the distance between the centroids of the top and bottom connecting surfaces. One can calculate them by doubling the averages of the ‘z’ coordinates calculated in Section 2. Therefore, the slant height for 1 pop-out is given by:
${\color[rgb]{1,0,0}(d_{\text{gold}})_{1-p/o}=2\left(\frac{A_{z}+B_{z}+C_{z}}{3}\right)=2\left(\frac{w+2w\sin\theta}{3}\right)=\frac{1}{3\varphi}\approx 0.206.}$ (35)
Similarly, for 2 pop-out:
${\color[rgb]{1,0,0}(d_{\text{gold}})_{2-p/o}=2\left(\frac{A_{z}+B_{z}+C_{z}}{3}\right)=2\left(\frac{w\sin\theta+2w}{3}\right)=\frac{2}{3\varphi}\approx 0.412.}$ (36)
As the 3 pop-out configuration (111) is fully opened, the total height is $2w$, more specifically,
${\color[rgb]{1,0,0}(d_{\text{gold}})_{3-p/o}=\tan(\beta)=\frac{1}{\varphi}\approx 0.618}$ (37)